# Quantum Thermoelectric Circuits: A Universal Approach

### v2

@@ Abstract
where the $\tilde{f}_{13}$ is a function of $\omega_{j},T_{j},J_{ij},T$, and $\gamma_{k}$ (for $i=1,2;j=1,2,3$ and $k=A,B$). Details of the proof are presented in [50]. Similarly, one can find out the various functions, such as $\tilde{f}_{12}$, and $\tilde{f}_{23}$, which can be used to identify the corresponding relation of the quantum thermal currents $\mathcal{J}_{12},\mathcal{J}_{21},\mathcal{J}_{23}$, and $\mathcal{J}_{32}$ with the quantum thermal potentials $V_{12}$ and $V_{23}$.
We use the above framework to develop the quantum thermal equivalent of an electric Wheatstone bridge and a quantum thermal adder circuit. Also, a diode and a quantum thermal transistor, from this perspective, are discussed in [50].
Quantum thermal Wheatstone bridge.— The classical Wheatstone bridge is a device primarily used to determine an unknown resistance. In a circuit of four resistances, one is a tunable resistance, and the other is the unknown, whose value is determined when the ratio between them becomes equal to that of the known resistances. In the case of the quantum Wheatstone bridge, the proposed device would determine an unknown interaction strength of a Hamiltonian. A pictorial comparison between a classical Wheatstone bridge and a quantum one is depicted in [51], wherein a model of quantum Wheatstone bridge was proposed, with the balance condition determining the unknown coupling strength obtained by observing a drop in population of Bell states caused by controlling a tunable coupling. On the other hand, our approach to finding the balance condition is via manipulating the heat current through the circuit, in tune with the actual classical Wheatstone bridge.
The Hamiltonian for the quantum Wheatstone bridge is given by $H_{S}^{W}=\frac{1}{2}\sum_{i=1}^{4}\omega_{i}\sigma^{z}_{i}+\sum_{l,k=1,l<k}^{4}J_{lk}\left(\sigma^{x}_{l}\sigma^{x}_{k}+\sigma^{y}_{l}\sigma^{y}_{k}\right)$, where $J_{12}=0$ as the qubits 1 and 2 are non-interacting. A schematic diagram of the quantum thermal Wheatstone bridge is shown in Fig. 5(a). This is similar to the setup of a Wheatstone bridge of an electrical circuit, where an unknown resistance can be determined using a specific balance condition when the current between nodes 3 and 4 is zero. Qubits 1 and 2 are impacted by the baths $A$ at zero temperature and $B$ at temperature $T,$ respectively. The evolution of the system is dictated by the quantum master equation of the form given in Eq. 1 by replacing $H_{S}$ with $H_{S}^{W}$, where the steady-state $\rho^{SS}_{W}$ can be obtained using $\mathcal{L}\left(\rho^{SS}_{W}\right)=0$. The balance condition for the Wheatstone bridge is that the quantum heat current $\mathcal{J}_{34}$ between qubits 3 and 4 should be zero. This happens when couplings follow the relation: $J_{13}=J_{14}$ and $J_{23}=J_{24}$. Note that $J_{13}$ need not be equal to $J_{23}$.
The phenomenon of the quantum Wheatstone Bridge is shown in Fig.  6. It can be observed that for the condition $J_{13}=J_{14}$, the red markers, depicting the current between qubits 3 and 4, are always zero. Further, green markers show the balance condition derived in [51], which is consistent at low temperatures; however, at high temperatures, the green markers shift slightly above the zero line, deviating from the balance condition.
@@
Quantum thermal adder circuit.– We now move on to discuss the quantum adder circuit, Fig. 5(b). This thermal circuit is motivated by the DC operational amplifier’s adder circuit, where the output voltage of the device is the sum of all the input voltages. Here, the qubits $1,2,...,n$ interact with the baths $I,II,...,N$, at temperatures $T_{I},T_{II},...,T_{N}$, respectively. The effective temperature of the qubits $1,2,...,n$ are $T_{1},T_{2},...T_{n}$, respectively. The thermal potential difference between the qubit 1 and bath $I$ is $V_{I1}=T_{I}-T_{1}$ and similarly for the other qubits. The qubits $1,2,...,n$ also interact with the qubit $\alpha$, which is connected to a zero temperature bath $T_{A}$ and the thermal potential difference between qubit $\alpha$ and bath $A$ is $V_{\alpha A}=T_{\alpha}-T_{A}=T_{\alpha}$, where $T_{\alpha}$ is the effective temperature of qubit $\alpha$. The Hamiltonian of the system is given by $H_{add}=\frac{\omega_{\alpha}}{2}\sigma^{z}_{\alpha}+\sum_{i=1}^{n}\left[\frac{\omega_{i}}{2}\sigma^{z}_{i}+J_{i\alpha}\left(\sigma^{x}_{\alpha}\sigma^{x}_{i}+\sigma^{y}_{\alpha}\sigma^{y}_{i}\right)\right]$. From Kirchhoff’s thermal current law, taking qubit $\alpha$ as a node, we can write $\mathcal{J}_{A\alpha}=\sum_{k=1}^{n}\mathcal{J}_{k\alpha}$ and using the relation $\mathcal{J}_{\alpha k}/\omega_{k}=-\mathcal{J}_{k\alpha}/\omega_{\alpha}$ (for $k=1,2,..,n$) from quantum thermal transformer, we can write $\mathcal{J}_{A\alpha}=-\sum_{k=1}^{n}\frac{\omega_{\alpha}}{\omega_{k}}\mathcal{J}_{\alpha k}$. But $\mathcal{J}_{\alpha k}=\mathcal{J}_{jk}$ (for $(j,k)=(I,1),(II,2),...(N,n)$); therefore, we get a relation between the input thermal currents $\mathcal{J}_{jk}$ and the output thermal current $\mathcal{J}_{A\alpha}$ as
@@
which for $\omega_{k}=\omega_{\alpha}\forall k$ results into $\mathcal{J}_{A\alpha}=-\left(\mathcal{J}_{I1}+\mathcal{J}_{II2}+...+\mathcal{J}_{Nn}\right)$. The above equation exactly mimics the relation between the input voltage and the output voltage of an adder circuit. As seen before, the currents $\mathcal{J}_{jk}$ and $\mathcal{J}_{A\alpha}$ are related to the effective temperature of the corresponding qubits $T_{k}$ via $\mathcal{J}_{jk}=\gamma_{j}\omega_{k}\left[-1+\coth\left(\frac{\omega_{k}}{2T_{j}}\right)\tanh\left(\frac{\omega_{1}}{2T_{k}}\right)\right],$ for $j=A,I,II,...N$ and $k=\alpha,1,2,...n$. Taking $\gamma_{A}=\gamma_{I}=...=\gamma_{N}$, $\omega_{1}=\omega_{2}=...=\omega_{n}=\omega$, $J_{1\alpha}=J_{2\alpha}=...=J_{n\alpha}$ and $T_{I}=T_{II}=...=T_{N}=T$, we get a simplified relation between the effective temperatures ($T_{x}$ for all qubits) of the qubits $1,2,...,n$ and the qubit $\alpha$ ($T_{\alpha}$), i.e.,
@@
Immediately, we observe that at lower temperatures $T$, the factor $\frac{2}{1+e^{\omega_{\alpha}/T_{\alpha}}}\rightarrow 0$. Interestingly, for this scenario, for higher values of $T$ and lower values of $\omega_{\alpha}$, the potential difference between bath $\mathcal{B}$ and qubits $V_{\mathcal{B}x}=T-T_{x}$ becomes independent of $T_{\alpha}$ and reduces to $V_{\mathcal{B}x}\approx\frac{T}{N+1}$. Remarkably, under certain parameter regimes, we get the condition that the sum of input thermal potentials becomes approximately equal to the output potential. For example, in the case of two qubits 1 and 2 interacting with qubit $\alpha$, keeping $\omega_{1}=\omega_{2}=\omega$, $J_{1\alpha}=J_{2\alpha}$, $T_{I}=T_{II}=T$, and $\gamma_{I}=\gamma_{II}=\gamma$, we get a condition on $\omega_{\alpha}$ ($\omega_{\alpha}=1.618N\omega$), and $\gamma_{A}$ ($\gamma_{A}=N\gamma$), such that $V_{I1}+V_{II2}=V_{\alpha A}$ with high accuracy in the high-temperature ($T$) regime. Similarly, for a higher number of qubits, $\omega_{\alpha}$ and $\gamma_{\alpha}$ can be tuned appropriately to get $V_{I1}+V_{II2}+...+V_{Nn}=V_{\alpha A}$, in general.
@@
Conclusion.— In this paper, we have presented a comprehensive schematic for a quantum thermal circuit theory, paving a possible roadmap for the construction of quantum thermal integrated circuits, important ingredients of which are quantum thermal diodes and transistors. We prove the existence of Kirchhoff-like voltage and current laws, with interaction strength playing the role of circuit resistance. We further demonstrate how heat currents can be manipulated in these quantum circuits to construct a quantum step transformer. Continuing this line of study, we demonstrate a Quantum Wheatstone bridge, where unknown Hamiltonian strengths can be determined by the profiling of heat currents, which could have a plethora of applications in process tomography, metrology, and quantum sensing. A model of a quantum thermal adder circuit for temperature gradients was also constructed, creating further opportunities to design thermal operational amplifiers and other integrated circuits. The framework of the quantum thermoelectric circuit developed here can be adapted to experimental realization. Thus, for example, progress has been made towards the diode architecture with pentamethyl-disilane in an NMR register [52]. Along similar lines, experimental demonstration of other quantum thermoelectric circuits discussed here could be envisaged. It is thus sufficient to say that the present work offers the fundamentals of a comprehensive quantum thermoelectric circuit theory, paving the way for the experimental demonstration of quantum thermoelectric integrated circuits.
@@
Acknowledgements.— SB2 acknowledges the support from the Ministry of Electronics and Information Technology, Government of India, under Grant No. 4(3)/2024-ITEA. SB1 acknowledges useful discussions with T.S. Mahesh on the possible experimental realizations of quantum thermal devices.
@@
Supplemental Material for “Quantum Thermoelectric Circuits: A Universal Approach”
@@ I Steady-state of quantum thermal resistor
$H_{S}$ $=H_{1}+H_{2}+H_{12}=\frac{\omega_{1}}{2}\sigma_{1}^{z}+\frac{\omega_{2}}{2}\sigma_{2}^{z}+J_{12}\left(\sigma^{x}_{1}\sigma^{x}_{2}+\sigma^{y}_{1}\sigma^{y}_{2}\right),$ (10)
where $\omega_{1}$ and $\omega_{2}$ are the transition frequencies for qubits 1 and 2, respectively, and $\sigma^{k}$’s (for $k=x,y,z$) are the Pauli spin matrices. $J_{12}$ is the interaction strength between qubits 1 and 2. To discuss the thermal transport between the two qubits and their respective baths, we find out the steady state of the system. The dynamics of the system under the Born-Markov and the secular approximations are given by the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation of the form.
@@ II Details on the forms of heat currents and proof of quantum thermal version of Kirchhoff’s current law
Consider a general system of $n$ qubits (labeled $1$, $2$, $3$, …, $n$), where each qubit is coupled to all other qubits. Further, each qubit is weakly coupled to its respective bosonic thermal bath (labeled $I$, $II$, $III$, …, $N$). The system’s Hamiltonian in this setup is given by
@@
where $H_{k}=\frac{\omega_{k}}{2}\sigma^{z}_{k}$ and $H_{lk}=J_{lk}\left(\sigma^{x}_{l}\sigma^{x}_{k}+\sigma^{y}_{l}\sigma^{y}_{k}\right)$ is the qubit-qubit interaction Hamiltonian with $J_{lk}$ being the coupling strength. Under the Born-Markov and rotating wave approximations, the dynamics of the system (depicted by $\rho$) is given by the GKSL master equation
@@
where $\mathcal{D}_{jk}(\rho)=\gamma_{j}(\tilde{N}_{jk}+1)\left(\sigma^{-}_{k}\rho\sigma^{+}_{k}-\frac{1}{2}\left\{\sigma^{+}_{k}\sigma^{-}_{k},\rho\right\}\right)+\gamma_{j}\tilde{N}_{jk}\left(\sigma^{+}_{k}\rho\sigma^{-}_{k}-\frac{1}{2}\left\{\sigma^{-}_{k}\sigma^{+}_{k},\rho\right\}\right)$, and $\tilde{N}_{jk}=\frac{1}{e^{\beta_{j}\omega_{k}}-1}$, with $\beta_{j}=T_{j}^{-1}$ for $(j,k)=(I,1),(II,2),...,(N,n)$. Let us pick a qubit $m$ (Hamiltonian $H_{m}$) from the $n$ qubits and consider it to be a node. This qubit $m$ would have quantum heat currents from all the sources, that is, from all the other qubits that interact with it as well as from the bosonic bath it is in interaction with (let the corresponding bath be $M$). The time derivative of the expectation value of $H_{m}$ is
@@
where the left-hand side of the above equation denotes the rate of change in the energy of the qubit $m$. On the right-hand side (RHS), the first term is the power, and the second term denotes the net heat current directed towards qubit $m$. Note that, in the present scenario, the Hamiltonian $H_{m}$ is constant; therefore, the first term on the RHS of the above equation becomes zero. Thus, $\frac{d}{dt}{\rm Tr}\left[H_{m}\rho\right]={\rm Tr}\left[H_{m}\frac{d\rho}{dt}\right]$. Now, we multiply Eq. (15) with $H_{m}$ as
@@
Taking trace on both sides in the above equation, we get
@@
The trace ${\rm Tr}\left[H_{m}\mathcal{D}_{jk}(\rho)\right]$ for $(j,k)=(I,1),(II,2),...,(N,n)$ in the above equation can be expanded as
${\rm Tr}\left[H_{m}\mathcal{D}_{jk}(\rho)\right]$ $={\rm Tr}\left[\gamma_{j}(\tilde{N}_{jk}+1)H_{m}\left(\sigma^{-}_{k}\rho\sigma^{+}_{k}-\frac{1}{2}\left\{\sigma^{+}_{k}\sigma^{-}_{k},\rho\right\}\right)\right]$
$+{\rm Tr}\left[\gamma_{j}\tilde{N}_{jk}H_{m}\left(\sigma^{+}_{k}\rho\sigma^{-}_{k}-\frac{1}{2}\left\{\sigma^{-}_{k}\sigma^{+}_{k},\rho\right\}\right)\right],$ (19)
which upon substituting $H_{m}=\frac{\omega_{m}}{2}\sigma^{z}_{m}$ becomes
${\rm Tr}\left[H_{m}\mathcal{D}_{jk}(\rho)\right]$ $=\gamma_{j}(\tilde{N}_{jk}+1)\frac{\omega_{m}}{2}{\rm Tr}\left[\left(\sigma^{z}_{m}\sigma^{-}_{k}\rho\sigma^{+}_{k}\right.\right.$
$-\left.\left.\frac{1}{2}\left\{\sigma^{z}_{m}\sigma^{+}_{k}\sigma^{-}_{k}\rho+\sigma^{z}_{m}\rho\sigma^{+}_{k}\sigma^{-}_{k}\right\}\right)\right]$
$+\gamma_{j}\tilde{N}_{jk}\frac{\omega_{m}}{2}{\rm Tr}\left[\left(\sigma^{z}_{m}\sigma^{+}_{k}\rho\sigma^{-}_{k}\right.\right.$
$-\left.\left.\frac{1}{2}\left\{\sigma^{z}_{m}\sigma^{-}_{k}\sigma^{+}_{k}\rho+\sigma^{z}_{m}\rho\sigma^{-}_{k}\sigma^{+}_{k}\right\}\right)\right].$ (20)
It can be easily verified from the above equation that ${\rm Tr}\left[H_{m}\mathcal{D}_{jk}(\rho)\right]=0$ for all $j,k$ except when $j=M$ and $k=m$. Therefore, Eq. (18) reduces to
@@
Substituting the above equation in Eq. (21), we get the net current directed towards qubit $m$ as
@@
It can be observed from the above equation that the heat current $\mathcal{J}_{lm}$ directed towards qubit $m$ from any qubit $l$ is given by
@@
and the heat current $\mathcal{J}_{Mm}$ from bath $M$ (attached to qubit $m$) to qubit $m$ is given by
@@
Further, in the steady state condition, the net current directed towards qubit $m$, Eq. (24), becomes zero, that is, ${\rm Tr}\left[H_{m}\frac{d\rho^{SS}}{dt}\right]=0$. This implies,

### v3

@@ Abstract
where the $\tilde{f}_{13}$ is a function of $\omega_{j},T_{j},J_{ij},T$, and $\gamma_{k}$ (for $i=1,2;j=1,2,3$ and $k=A,B$). Similarly, one can find out the various functions, such as $\tilde{f}_{12}$, and $\tilde{f}_{23}$, which can be used to identify the corresponding relation of the quantum thermal currents $\mathcal{J}_{12},\mathcal{J}_{21},\mathcal{J}_{23}$, and $\mathcal{J}_{32}$ with the quantum thermal potentials $V_{12}$ and $V_{23}$.
We use the above framework to develop novel circuits, to which we coin the names quantum thermal super Wheatstone bridge and quantum thermal adder. Further, thermal devices discussed in recent times, for example, a Wheatstone bridge, diode, and quantum thermal transistor, are presented using our framework in [54].
Quantum thermal super Wheatstone bridge—We discuss the circuit for a quantum thermal super Wheatstone Bridge, Fig. 5, with Hamiltonian $H_{S}^{SW}=\sum_{j=1}^{7}\frac{\omega_{j}}{2}\sigma^{z}_{j}+H_{15}+H_{16}+H_{23}+H_{24}+H_{34}+H_{35}+H_{37}+H_{46}+H_{47}+H_{56}+H_{57}+H_{67}$, where $H_{jk}=J_{jk}\left(\sigma^{x}_{j}\sigma^{x}_{k}+\sigma^{y}_{j}\sigma^{y}_{k}\right)$. The dynamics of the system is governed by Eq. (1), by replacing $H_{S}$ with $H_{S}^{SW}$. The geometry of this circuit makes it very interesting. We discuss three important facets of this circuit here. In condition 1, the circuit is made symmetrical, such that all the $\omega_{j}$’s as well as all the $J_{jk}$’s are the same ($J$). In this condition, we observe that the currents $\mathcal{J}_{34}$ and $\mathcal{J}_{56}$ become zero, see Fig. 6(a).
Figure 6: Variation of thermal currents $\mathcal{J}_{jk}$ with interaction strengths $J_{jk}$. (a), (b) and (c) depict conditions 1, 2, and 3 of the super Wheatstone bridge, mentioned in the text below. In all the plots, unless conditioned for the balanced state of the bridge, the values of $J_{jk}$’s are of the order of $\omega_{j}$’s, which are randomly taken between 1 and 2. Further, $T=10,\gamma_{A}=0.1$, and $\gamma_{B}=0.05$.
In condition 2, we explore the scenario where the center qubit 7 is not affected by the thermal transport around it. To do so, we allow the input and output of heat from single connections to the square made by qubits 3, 4, 5, and 6, thereby making $J_{23}=J_{61}=0$. Further, we allow only two connections to qubit $7$, for example, with qubit 3 and 6, implying $J_{57}=J_{47}=0$. The balance condition is obtained when $J_{34}=J_{46}$ and $J_{35}=J_{56}$ ($J_{34}$ need not be equal to $J_{35}$), together with $\omega_{3}=\omega_{6}=\omega_{7}$. In this condition, see Fig. 6(b), the currents $\mathcal{J}_{37}=\mathcal{J}_{67}=0$, making the qubit 7 unaffected by the thermal transport around it. Interestingly, if we increase the number of qubits between qubits 3 and 6, keeping their $\omega_{i}$’s equal, they will also remain unaffected by the thermal transport around them [54]. Condition 3 of the super Wheatstone bridge is identical to the Wheatstone bridge condition [54]. We can form many Wheatstone bridges inside this super Wheatstone bridge; for example, see the black dashed line between qubits 1, 5, 7, and 6. Here, we take $J_{35}=J_{37}=J_{46}=0$ to form a particular Wheatstone bridge. The balance condition is achieved when $J_{15}=J_{16}$ and $J_{57}=J_{67}$ ($J_{15}$ need not be equal to $J_{57}$); and then the current between qubits 5 and 6, $\mathcal{J}_{56}$, becomes zero, see Fig. 6(c), a classic Wheatstone bridge condition.
Quantum thermal adder circuit—We now discuss the quantum thermal adder circuit, Fig. 7.
Figure 7: A schematic diagram of the quantum thermal adder circuit.
@@
This thermal circuit is motivated by the DC operational amplifier’s adder circuit, where the output voltage of the device is the sum of all the input voltages. Here, the qubits $1,2,...,n$ interact with the baths $I,II,...,N$, at temperatures $T_{I},T_{II},...,T_{N}$, respectively. The effective temperatures of the qubits $1,2,...,n$ are $T_{1},T_{2},...,T_{n}$, respectively. The thermal potential difference between qubit 1 and bath $I$ is $V_{I1}=T_{I}-T_{1}$ and similarly for the other qubits. Qubits $1,2,...,n$ also interact with qubit $\alpha$ connected to a zero temperature bath $T_{A}$ and the thermal potential difference between qubit $\alpha$ and bath $A$ is $V_{\alpha A}=T_{\alpha}-T_{A}=T_{\alpha}$, where $T_{\alpha}$ is the effective temperature of qubit $\alpha$. The Hamiltonian of the system is $H_{add}=\frac{\omega_{\alpha}}{2}\sigma^{z}_{\alpha}+\sum_{i=1}^{n}\left[\frac{\omega_{i}}{2}\sigma^{z}_{i}+J_{i\alpha}\left(\sigma^{x}_{\alpha}\sigma^{x}_{i}+\sigma^{y}_{\alpha}\sigma^{y}_{i}\right)\right]$. Using the thermal current law, $\mathcal{J}_{A\alpha}=\sum_{k=1}^{n}\mathcal{J}_{k\alpha}$ and Eq. (5), we can write $\mathcal{J}_{A\alpha}=-\sum_{k=1}^{n}\frac{\omega_{\alpha}}{\omega_{k}}\mathcal{J}_{\alpha k}$. But $\mathcal{J}_{\alpha k}=\mathcal{J}_{jk}$ (for $(j,k)=(I,1),(II,2),...(N,n)$), implying the following relation
@@
which for $\omega_{k}=\omega_{\alpha}\forall k$ results into $\mathcal{J}_{A\alpha}=-\left(\mathcal{J}_{I1}+\mathcal{J}_{II2}+...+\mathcal{J}_{Nn}\right)$. The above equation exactly mimics the relation between the input and output voltages of an adder circuit. The currents $\mathcal{J}_{jk}$ and $\mathcal{J}_{A\alpha}$ are related to the effective temperature of the corresponding qubits $T_{k}$ via $\mathcal{J}_{jk}=\gamma_{j}\omega_{k}\left[-1+\coth\left(\frac{\omega_{k}}{2T_{j}}\right)\tanh\left(\frac{\omega_{k}}{2T_{k}}\right)\right],$ for $j=A,I,II,...N$ and $k=\alpha,1,2,...n$. Taking $\gamma_{A}=\gamma_{I}=...=\gamma_{N}$, $\omega_{1}=\omega_{2}=...=\omega_{n}=\omega$, $J_{1\alpha}=J_{2\alpha}=...=J_{n\alpha}$ and $T_{I}=T_{II}=...=T_{N}=T$, we get a simplified relation between the effective temperatures $T_{x}$ for $x=1,2,...,n$ and the qubit $\alpha$ ($T_{\alpha}$)
@@
We observe that at lower temperatures $T$, $\frac{2}{1+e^{\omega_{\alpha}/T_{\alpha}}}\rightarrow 0$. Interestingly, for higher values of $T$ and lower values of $\omega_{\alpha}$, the potential difference between bath $\mathcal{B}$ and qubits $V_{\mathcal{B}x}=T-T_{x}$ becomes independent of $T_{\alpha}$ and reduces to $V_{\mathcal{B}x}\approx\frac{T}{N+1}$. Remarkably, under certain parameter regimes, we get the condition that the sum of input thermal potentials becomes approximately equal to the output potential. For example, for two qubits 1 and 2 interacting with qubit $\alpha$, keeping $\omega_{1}=\omega_{2}=\omega$, $J_{1\alpha}=J_{2\alpha}$, $T_{I}=T_{II}=T$, and $\gamma_{I}=\gamma_{II}=\gamma$, we get a condition on $\omega_{\alpha}$ ($\omega_{\alpha}=1.618N\omega$), and $\gamma_{A}$ ($\gamma_{A}=N\gamma$), such that $V_{I1}+V_{II2}=V_{\alpha A}$ with high accuracy in the high-temperature ($T$) regime. Similarly, for a higher number of qubits, $\omega_{\alpha}$ and $\gamma_{\alpha}$ can be tuned appropriately to get $V_{I1}+V_{II2}+...+V_{Nn}=V_{\alpha A}$, in general.
@@
Conclusion—We have presented a comprehensive schematic for a quantum thermal circuit theory, paving a roadmap to construct quantum thermal integrated circuits, the important ingredients of which are quantum thermal diodes and transistors. We prove the existence of Kirchhoff-like voltage and current laws, with interaction strength playing the role of circuit resistance. We further demonstrate how heat currents can be manipulated to construct a quantum step transformer. Using this, we develop a quantum super Wheatstone bridge with many interesting balance conditions, also containing the standard quantum Wheatstone bridge, where unknown Hamiltonian strengths can be determined by the profiling of heat currents. This circuit can have a plethora of applications in process tomography, metrology, and quantum sensing. A model of a quantum thermal adder circuit for temperature gradients was also constructed, creating further opportunities to design thermal operational amplifiers and other integrated circuits. We also bring out the interrelation between the heat current developed here and the established forms of current used in the literature. Thus, for example, the spin current, which can be experimentally measured [60, 61, 53], provides a handle for experimental determination of the heat current. The framework of the circuits developed here can be adapted to experimental realization. Thus, for example, progress has been made towards the diode architecture with pentamethyl-disilane in an NMR register [62]. Also, an experimental study of thermodynamic uncertainty relation in the two-qubit NMR spin system, sodium fluorophosphate, was performed in [63], and an experimental setup of a quantum thermal transistor using a superconducting circuit was proposed in [64]. Along similar lines, experimental demonstration of other quantum thermal circuits discussed here could be envisaged. Our analysis opens up several research avenues, for example, studies in thermal conductivity, balance conditions of the super Wheatstone bridge, and other analogs of electric circuits. It is thus sufficient to say that the present work offers the fundamentals of a comprehensive theory of quantum thermal analogs of electric circuits, paving the way for the corresponding integrated circuits.
@@
Acknowledgments—S. Bhattacharya acknowledges the support from the Ministry of Electronics and Information Technology, Government of India, under Grant No. 4(3)/2024-ITEA. S. Banerjee acknowledges useful discussions with T.S. Mahesh on the possible experimental realizations of quantum thermal devices.
@@
Supplemental Material for “Quantum Thermal Analogs of Electric Circuits: A Universal Approach”
@@ I Steady-state of quantum thermal resistor
$H_{S}$ $=H_{1}+H_{2}+H_{12}$
$=\frac{\omega_{1}}{2}\sigma_{1}^{z}+\frac{\omega_{2}}{2}\sigma_{2}^{z}+J_{12}\left(\sigma^{x}_{1}\sigma^{x}_{2}+\sigma^{y}_{1}\sigma^{y}_{2}\right),$ (10)
where $\omega_{1}$ and $\omega_{2}$ are the transition frequencies for qubits 1 and 2, respectively, and $\sigma^{k}$’s (for $k=x,y,z$) are the Pauli spin matrices. $J_{12}$ is the interaction strength between qubits 1 and 2. To discuss the thermal transport between the two qubits and their respective baths, we find the steady state of the system. The dynamics of the system under the Born-Markov and the secular approximations are given by the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation of the form
@@ II Details on the forms of heat currents and proof of quantum thermal version of Kirchhoff’s current law
Consider a general system of $n$ nodes (labeled $1$, $2$, $3$, …, $n$), where each node is coupled to all other nodes. Further, each node is weakly coupled to its respective bosonic thermal bath (labeled $I$, $II$, $III$, …, $N$). The system’s Hamiltonian in this setup is given by
@@
where $H_{k}$ is the Hamiltonian of the node (here, it is $H_{k}=\frac{\omega_{k}}{2}\sigma^{z}_{k}$) and $H_{lk}$ is the coupling Hamiltonian between the nodes (here, we take $H_{lk}=J_{lk}\left(\sigma^{x}_{l}\sigma^{x}_{k}+\sigma^{y}_{l}\sigma^{y}_{k}\right)$ with $J_{lk}$ being the coupling strength). Under the Born-Markov and rotating wave approximations, the dynamics of the system (depicted by $\rho$) is given by the GKSL master equation
@@
where $\mathcal{D}_{jk}(\rho)=\gamma_{j}(\tilde{N}_{jk}+1)\left(L_{k}\rho L_{k}^{\dagger}-\frac{1}{2}\left\{L_{k}^{\dagger}L_{k},\rho\right\}\right)+\gamma_{j}\tilde{N}_{jk}\left(L_{k}^{\dagger}\rho L_{k}-\frac{1}{2}\left\{L_{k}L_{k}^{\dagger},\rho\right\}\right)$, and $\tilde{N}_{jk}=\frac{1}{e^{\beta_{j}\omega_{k}}-1}$, with $\beta_{j}=T_{j}^{-1}$ for $(j,k)=(I,1),(II,2),...,(N,n)$. Further, $L_{k}$ are the Lindblad jump operators, which in the case of qubits are $\sigma^{-}_{k}$. Let us pick a node $m$ (Hamiltonian $H_{m}$) from the $n$ nodes. This node $m$ would have quantum heat currents from all the sources, that is, from all the other nodes that interact with it, as well as from the bosonic bath it is in interaction with (let the corresponding bath be $M$). The time derivative of the expectation value of $H_{m}$ is
@@
where the left-hand side of the above equation denotes the rate of change in the energy of the node $m$. On the right-hand side (RHS), the first term is the power, and the second term denotes the net heat current directed towards node $m$. Note that, in the present scenario, the Hamiltonian $H_{m}$ is constant; therefore, the first term on the RHS of the above equation becomes zero. Thus, $\frac{d}{dt}{\rm Tr}\left[H_{m}\rho\right]={\rm Tr}\left[H_{m}\frac{d\rho}{dt}\right]$. Now, we multiply Eq. (15) with $H_{m}$ as
@@
Taking the trace on both sides in the above equation, we get
@@
It can be easily verified from the above equation that ${\rm Tr}\left[H_{m}\mathcal{D}_{jk}(\rho)\right]=0$ for $(j,k)=(I,1),(II,2),...,(N,n)$ except when $j=M$ and $k=m$. Therefore, Eq. (18) reduces to
@@
Substituting the above equation in Eq. (19), we get the net current directed towards node $m$ as
@@
It can be observed from the above equation that the heat current $\mathcal{J}_{lm}$ directed towards node $m$ from any other node $l$ is given by
@@
and the heat current $\mathcal{J}_{Mm}$ from bath $M$ (attached to node $m$) to node $m$ is given by
@@
Further, in the steady state condition, the net current directed towards node $m$, Eq. (22), becomes zero, that is, ${\rm Tr}\left[H_{m}\frac{d\rho^{SS}}{dt}\right]=0$. This implies,